# Machine learning applications in archaeological practices: a review

### v3

@@ 6.1 Glossary
  • CNN = Convolutional neural network
  • DGCNN = Dynamic graph convolutional neural network
  • FNN = Feedforward neural network
  • FR-CNN = Faster region based convolutional neural network
  • GAN = Generative adversarial network
@@
  • GSSL = Graph-based semi-supervised learning
@@
  • ISODATA = Iterative self-organising data analysis
  • k-MC = k-mean clustering
  • kNN = k-nearest neighbors
@@
  • MR-CNN = Mask region based convolutional neural network
@@
  • R-CNN = Region based convolutional neural network
  • RF = Random forest
@@
  • SBS-RFR = Sequential backward selection-random forest regression
  • SLIC = Simple linear iterative clustering
  • SVM = Support vector machine

### v5

@@ 6 Annexes
Table 4: List of algorithms present in the study cases reviewed, along with their abbreviations, number of uses in our corpus, and number of times they performed best in comparative evaluations.
Description Abbreviation Uses Best performance
Artificial Neural Network
Feedforward Neural Network FNN 23 4
Convolutional Neural Network CNN 14 1
Residual Neural Network ResNet 12 2
Mask Region-based Convolutional Neural Network MR-CNN 9 1
Faster Region-based Convolutional Neural Network FR-CNN 8 0
Visual Geometry Group VGG 8 2
U-Net U-Net 7 4
Inception Network INC 4 1
AlexNet AlexNet 3 0
RetinaNet RN 3 0
YOLO YOLO 3 0
DeepLabv3+ DL3 2 0
Semantic Segmentation Model SegNet 2 0
Adaptive deep learning for fine-grained image retrieval ADLFIG 1 0
Bidirectional Encoder Representations from Transformers (DNLM) BERT 1 0
BiGRU-Dual Attention BiGRU 1 0
BiLSTM (RNN/Recurrent Neural Network) BiLSTM 1 0
Dynamic Graph Convolutional Neural Network DGCNN 1 0
DenseNet201 DN201 1 0
Generative Adversarial Network GAN 1 0
Jason 2 JAS2 1 0
Neural Support Vector Machine NSVM 1 0
Region-based Convolutional Neural Network R-CNN 1 0
SimpleNet SimpleNet 1 0
Single Shot MultiBox Detector SSD 1 0
Bayesian Classifier
Naïve Bayes NB 11 0
Maximum Entropy MaxEnt 2 0
Decision Trees and Rule Induction
C5.0 C5.0 7 2
C4.5 C4.5 4 0
Decision Tree / Classification Tree DT 4 0
Conditional Inference Trees CTREE 2 0
Iterative Dichotomiser 3 ID3 2 0
Classification And Regression Tree CART 1 0
Fast and Frugal Tree FFT 1 0
Learning with Galois Lattice LEGAL 1 0
Representative Trees REPTree 1 0
Random Tree RT 1 0
Ensemble Learning
Random Forest RF 54 20
Adaptive Boost AdaBoost 2 0
Stochastic Gradient Boosting SGB 2 1
eXtreme Gradient Boosting XGB 2 1
Bootstrap Aggregating BAgg 1 0
Discrete SuperLearner dSL 1 0
Fast Random Forest FRF 1 0
Gradient Boosting Regression Tree GboostRT 1 0
LogitBoost LB 1 0
Quantile Random Forest qRF 1 0
Sequential Backward Selection–Random Forest Regression SBS-RFR 1 1
SMOTE Boost SMOTEBoost 1 0
SMOTE + Edited Nearest Neighbour Rule SMOTEENN 1 0
Super Learner SP 1 1
Viola-Jones Cascade Classifier VL-$C$ 1 0
Genetic Algorithm GA 1 0
Linear Classifier
Support Vector Machine SVM 26 2
Structured SVM SSVM 1 0
Nearest Neighbour Classifier
k-nearest neighbors kNN 19 1
Weighted k-nearest neighbors kkNN 3 0
Polynomial Classifier
Support Vector Machine with Radial Basis Function Kernel SVMr 7 1
Unsupervised Learning and Clustering
Affinity Propagation AF 1 0
Hierarchical Cluster-Based Peak Alignment CluPA 1 0
Databionic Swarm DBS 1 0
Expectation-Maximisation Clustering EMC 1 0
Graph-based Semi-Supervised Learning GSSL 1 1
Iterative Closest Point ICP 1 0
Iterative Self-Organizing Data Analysis ISODATA 1 0
Nearest Centroid NC 1 0
Simple Linear Iterative Clustering SLIC 1 0
Self-Organizing Map SOM 1 0
Tilburg Memory-Based Learning TiMBL 1 0
Time series clustering TSC 1 0
k-Mean Clustering k-MC 7 0
Table 4: (continued)